f is a function defined by : f(x) = x^2-3.

Find f(0), f(2) and f(4).


1
Expert's answer
2021-11-08T16:30:28-0500

Given that "f(x)=x^2-3"

Put "x=0," we get:

"f(0)=(0)^2-3=-3"

Put "x=2," we get:

"f(2)=(2)^2-3=1"

Put "x=4," we get:

"f(4)=(4)^2-3=13"